# Defending Against Weight-Poisoning Backdoor Attacks for Parameter-Efficient Fine-Tuning

### v2

@@ Appendix A Experimental Setting
We have selected five popular NLP models as victim: BERT-large (Kenton and Toutanova,, 2019), RoBERTa-large (Liu et al.,, 2019), LLaMA-7B (Touvron et al.,, 2023), Vicuna-7B (Zheng et al.,, 2023) and MPT-7B (Team,, 2023). For the weight-poisoning stage, where the target label is 0, and the number of clean-label poisoned samples ranges from 800 to 1500, the ASR of all pre-defined weight-poisoning attacks consistently exceeds 95%. We adopt the Adam optimizer to train the classification model. For LoRA, we set the rank $r$ to 8 and dropout to 0.1. In the case of Prompt-tuning, P-tuning v1, and P-tuning v2, we set the virtual token to {4, 5}, the encoder hidden size to {64, 128}, the learning rate to {2e-5, 2e-3} for different fine-tuning strategies, the batch size to {32, 8}, and the threshold $\gamma$ to {0.7, 0.75} for different models. We perform all experiments on NVIDIA RTX A6000 GPU with 48G memory. Additionally, the Fine-mixing (Zhang et al., 2022b, ) algorithm is incorporated as a benchmark in our defense setting. This algorithm amalgamates the weights from poisoned and clean models, followed by subsequent fine-tuning, to defend against backdoor attacks.
Appendix B More Experiments Results
@@
Model Sample 0.5-0.6 0.6-0.7 0.7-0.8 0.8-0.9 0.9-1.0
Victim Poison 3% 4% 13% 31% 49%
PSIM Clean 70% 28% 2% 0% 0%
PSIM Poison 0% 1% 11% 31% 56%
Table 12: The distribution results of confidence scores from victim and PSIM module. The dataset is SST-2. Victim stands for a victim model.
To further substantiate our conjecture and evaluate the universality of our proposed defensive strategies, we have undertaken tests in intricate classification scenarios utilizing the AG’s News dataset (Zhang et al.,, 2015), which is a multiclass classification. The empirical outcomes are delineated in Table 8. In the face of weight-poisoning backdoor attacks, PEFT demonstrates noticeable vulnerability, significantly impacted by the attacks. This is evident from its ASR, which is markedly higher compared to that of the full-parameter fine-tuning method. Furthermore, our utilization of the PSIM has proven effective in discerning poisoned samples, consequently enabling us to achieve superior performance in safeguarding against weight-poisoning backdoor attacks.
PSIM in more language models To further validate the security issues of the PEFT algorithm when facing weight-poisoning backdoor attacks and to assess the generalizability of the PSIM algorithm, we conduct experiments on the Vicuna-7B (Zheng et al.,, 2023) and MPT-7B (Team,, 2023) models. As Table 10 shows, the experimental results indicate that the PEFT method exhibits a higher attack success rate when subjected to weight-poisoning backdoor attacks, which further corroborates our hypothesis that the PEFT method is more susceptible to such attacks. Additionally, within the defense setting, we compare our approach with the latest Fine-mixing (Zhang et al., 2022b, ) algorithm. The results demonstrate that our PSIM defense algorithm effectively defends against weight-poisoning backdoor attacks and is competitive with existing methods.
@@
PSIM in instruction tuning Unlike traditional supervised learning, instruction tuning (Xu et al.,, 2023; Yan et al.,, 2023) may not require fine-tuning third-party models, thereby naturally avoiding the issue of "catastrophic forgetting" during the fine-tuning process. However, if the weights are poisoned, the pre-defined backdoor attack trigger easily induces the model to output target content. We attempt to design a backdoor attack based on weight-poisoning for instruction tuning while using our algorithm for defense. The results are shown in Table 11; when utilizing the weight-poisoning language model directly, the attack success rate is close to 100%. When facing the defense algorithm, our PSIM effectively identifies poisoned samples, defends against backdoor attacks, and ensures the model’s performance. For example, in the LLaMA model, our PSIM algorithm achieves 100% defense without affecting model performance.
Statistical analysis for confidence About the setting of $\gamma$, we determine it through human experience. Unlike traditional hyperparameters, the selection of $\gamma$ does not impact the training of the PSIM module. We analyze the confidence scores of model outputs in the setting of weight-poisoning backdoor attacks based on BadNet. Firstly, we fine-tune the poisoned weights on clean samples utilizing the PEFT algorithm. During the inference phase, when the input samples contain the trigger, the proportion of cases where the model’s confidence score for the target label exceeds 0.8 is 80%. We also compile the confidence scores outputted by the PSIM module, which include the confidence of clean and poisoned samples towards the target label. It is not hard to see that when the input to the PSIM module is a clean sample, its confidence score tends to be around 0.5, while if the input sample is poisoned, the confidence score outputted by the module concentrates above 0.8. Therefore, we choose to set this parameter through human experience.
Label Reset Rate In our algorithm, random label resetting is utilized for training the PSIM module, thereby enabling the distinction between clean and poisoned samples based on confidence scores. Consequently, we necessitate random resetting of all labels within the samples. Concurrently, to evaluate the influence of varying label reset rates on defensive performance and clean accuracy, we conducted ablation experiments. As indicated in Table 13, despite the stability of defensive performance, a continuous decrease in label reset rate severely impairs clean accuracy. Therefore, it is indispensable to reset the labels of all samples within the PSIM module.
Scenario 100% 80% 60% 40%
CA ASR CA ASR CA ASR CA ASR
Full-tuning 91.08 4.65 89.13 6.05 60.41 0 34.98 0
LoRA 90.02 7.92 87.70 15.93 58.92 0.33 33.72 0.33
Table 13: Results of different label reset rates
Defense Method Extra module Graphics Storage
WeDef (Jin et al.,, 2022) Yes Yes Yes
Fine-mixing (Zhang et al., 2022b, ) Yes Yes Yes
DARCy (Le et al.,, 2020) Yes Yes No
AttDef (Li et al., 2023b, ) Yes Yes Yes
PSIM Yes Yes Yes
Table 14: Comparison of graphics memory and storage consumption across different backdoor attack defense algorithms.
Communication cost In our defense algorithm, we design the PSIM module, which is trained based on the PEFT algorithm. For example, in the LLaMA-7B model, we use the prompt-tuning algorithm to train the PSIM module. The number of parameters during the training process of this module is 1,097,984, which accounts for 0.016% of the total parameters (6,608,449,792) of the LLaMA model. Therefore, we believe that the graphics memory consumption for training an additional PSIM module is extremely low. In addition, we only need storage space equivalent to that of LLaMA to store the PSIM module. Furthermore, compared with several existing defense algorithms in Table 14, we found that the PSIM module does not increase the occupancy of graphics memory or storage space compared to current methods. For instance, in the Fine-mixing (Jin et al.,, 2022) defense algorithm, an additional model that is free of backdoors is required to be mixed proportionally with the weight-poisoned language model. This algorithm also necessitates extra graphics memory and storage space. Similarly, in AttDef (Li et al., 2023b, ), a poison sample discriminator based on ELECTRA is trained, which likewise requires additional graphics memory and storage space.

### v3

@@ Appendix B Experimental Setting
We have selected five popular NLP models as victim: BERT-large Kenton and Toutanova (2019), RoBERTa-large Liu et al. (2019), LLaMA-7B Touvron et al. (2023), Vicuna-7B Zheng et al. (2023) and MPT-7B Team (2023). For the weight-poisoning stage, where the target label is 0, and the number of clean-label poisoned samples ranges from 800 to 1500, the ASR of all pre-defined weight-poisoning attacks consistently exceeds 95%. We adopt the Adam optimizer to train the classification model. For LoRA, we set the rank $r$ to 8 and dropout to 0.1. In the case of Prompt-tuning, P-tuning v1, and P-tuning v2, we set the virtual token to {4, 5}, the encoder hidden size to {64, 128}, the learning rate to range from 2e-5 to 2e-3 for different fine-tuning strategies, the batch size to {32, 8}, and the threshold $\gamma$ to {0.7, 0.75} for different models. We perform all experiments on NVIDIA RTX A6000 GPU with 48G memory. Additionally, the Fine-mixing Zhang et al. (2022b) algorithm is incorporated as a benchmark in our defense setting. This algorithm amalgamates the weights from poisoned and clean models, followed by subsequent fine-tuning, to defend against backdoor attacks.
Appendix C More Experiments Results
@@
To further substantiate our conjecture and evaluate the universality of our proposed defensive strategies, we have undertaken tests in intricate classification scenarios utilizing the AG’s News dataset Zhang et al. (2015), which is a multiclass classification. The empirical outcomes are delineated in Table 8. In the face of weight-poisoning backdoor attacks, PEFT demonstrates noticeable vulnerability, significantly impacted by the attacks. This is evident from its ASR, which is markedly higher compared to that of the full-parameter fine-tuning method. Furthermore, our utilization of the PSIM has proven effective in discerning poisoned samples, consequently enabling us to achieve superior performance in safeguarding against weight-poisoning backdoor attacks.
PSIM in more language models To further validate the security issues of the PEFT algorithm when facing weight-poisoning backdoor attacks and to assess the generalizability of the PSIM algorithm, we conduct experiments on the Vicuna-7B Zheng et al. (2023) and MPT-7B Team (2023) models. As Table 10 shows, the experimental results indicate that the PEFT method exhibits a higher attack success rate when subjected to weight-poisoning backdoor attacks, which further corroborates our hypothesis that the PEFT method is more susceptible to such attacks. Additionally, within the defense setting, we compare our approach with the latest Fine-mixing Zhang et al. (2022b) algorithm. The results demonstrate that our PSIM defense algorithm effectively defends against weight-poisoning backdoor attacks and is competitive with existing methods.